# Quantized Hall conductivity in monolayer 1T-MoS2

### v2

@@ IV Conclusions
In conclusion, our investigation of spin-valley-resolved Hall conductivity and thermoelectric effects offers valuable insights into the topological phase transitions in 1T-MoS2. Our results show coexistence of the chiral and helical spin textures in the system. Changing $\alpha$ is responsible for the sign reversal of the Berry curvature near the Dirac points. The interplay between the spin-resolved and valley-resolved Hall conductivity underscores the influence of the parameter $\alpha$ on the topological phase transition from the QSHI-to-BI phase. This transition is further corroborated by the evolution of topological invariants, such as $C_{s}$ and $C_{v}$, aligning with previous studies. Additionally, the calculations of the SNC, VNC, and TNC reveal the dominance of SNC in the QSHI phase, which diminishes as VNC takes precedence in the BI phase, reinforcing the trends observed in Hall conductivity.

### v5

@@ IV Conclusions
In conclusion, this study uncover the topological phases and properties of the 1T-MoS2 theoretically. The analysis of spin textures reveals hallmark features such as spin-momentum locking, anisotropy in the band structure, and the presence of Dirac points, underscoring their potential for spintronic applications. The Berry curvature and Hall conductivity analyses demonstrate that the system undergoes a topological phase transition from a QSHI-to-BI phase as the parameter $\alpha$ is varied, with distinct spin-resolved and valley-resolved contributions to the Hall conductivity. Furthermore, the behavior of the spin and valley Nernst coefficients reveals nuanced energy-dependent transport phenomena, where significant contributions from spin and valley effects largely cancel each other, resulting in a zero total Nernst coefficient.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Conflict of interest
The authors have no conflicts to disclose.